se the following data and determine the se.

x y (y − ŷ)
5 45 3.798
7 38 -0.214
11 32 -0.236
12 24 -6.742
21 22 4.708
25 10 -1.315
Answer #1

from above:

x y (y − ŷ) (y − ŷ)^2
5 45 3.798 14.4248
7 38 -0.214 0.045796
11 32 -0.236 0.055696
12 24 -6.742 45.45456
21 22 4.708 22.16526
25 10 -1.315 1.729225
total 83.87535

SSE =83.87535

therefore se =sqrT(SSE/(n-2))=sqrt(83.87535/4)=4.5792